OpenStudy (anonymous):

Find S. A. 45° B. 80° C. 90° D. 100°

OpenStudy (anonymous):

what is the picture

OpenStudy (anonymous):

all the angles added together equal 360

OpenStudy (mrnood):

Intereior anglea of quadrilateral add up to 360 deg add all the given angles and solve for x - and then for S

OpenStudy (anonymous):

ITS 225 BUT I STILL HAVE TO SOLVE FOR X IN ORDER TO GET MEASURE S

OpenStudy (mrnood):

yes you do write down the sum of th eangles (including the x's) R+S+T+U =? = 360

OpenStudy (anonymous):

wat is 225??

OpenStudy (anonymous):

145+80=225

OpenStudy (mrnood):

now add on the 2 other angles including the x's

OpenStudy (anonymous):

if you do... 360=145+80+x+2x combine like terms and solve for x you will get your answer

OpenStudy (anonymous):

2X+X=3X BUT IF I DIVIDE 225 BY 3 IT GNNA GIVE ME 75

OpenStudy (mrnood):

just slow down - and read the help you are being given do as I suggested or as @Kitty_Kat has shown oyu...

OpenStudy (anonymous):

360-145-80=3x

OpenStudy (anonymous):

360-145-80=135

OpenStudy (anonymous):

135 DIVIDED BY 3X WOULD GIVE ME 45

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

solve for s

OpenStudy (anonymous):

145+45+80=90